# FLUID: Training-Free Face De-identification via Latent Identity Substitution

### v1

@@ Conclusion
8 steps 0.858 0.999 0.579 0.873 0.610 0.543 0.799
12 steps 0.762 1.000 0.586 0.843 0.630 0.617 0.782
16 steps 0.739 1.000 0.620 0.883 0.646 0.638 0.786
20 steps 0.651 0.999 0.634 0.876 0.681 0.641 0.755
24 steps 0.624 1.000 0.616 0.870 0.679 0.652 0.741
Table 3: Quantitative results on the number of denoising steps used for optimization.
@@
$\text{All}=1.0$ 0.686 1.000 0.602 0.844 0.615 0.670 0.759
$\lambda_{\text{id}}=0.5$ 0.659 1.000 0.596 0.895 0.666 0.668 0.756
$\lambda_{\text{id}}=1.5$ 0.775 0.999 0.532 0.876 0.648 0.666 0.788
$\lambda_{\text{attr}}=0.5$ 0.751 1.000 0.493 0.830 0.624 0.657 0.765
$\lambda_{\text{attr}}=1.5$ 0.670 1.000 0.593 0.888 0.639 0.656 0.756
$\lambda_{\text{mask}}=0.5$ 0.737 1.000 0.620 0.883 0.646 0.638 0.786
$\lambda_{\text{mask}}=1.5$ 0.679 1.000 0.541 0.880 0.646 0.675 0.754
Table 4: Quantitative results of ablation study on each loss function weight, with updated metrics and nested harmonic‐mean “Overall.” Other weights not mentioned for each case are set to the default value of $1.0$.
[…]
100 0.569 1.000 0.653 0.915 0.696 0.665 0.723
200 0.641 1.000 0.659 0.902 0.677 0.654 0.756
300 0.674 1.000 0.643 0.904 0.671 0.655 0.768
400 0.692 0.999 0.630 0.893 0.668 0.646 0.772
500 0.695 1.000 0.663 0.899 0.665 0.632 0.776
600 0.710 1.000 0.651 0.880 0.646 0.636 0.778
700 0.716 1.000 0.620 0.872 0.627 0.635 0.773
800 0.721 1.000 0.634 0.870 0.624 0.630 0.776
900 0.722 1.000 0.620 0.877 0.640 0.637 0.778
1000 0.739 1.000 0.620 0.883 0.646 0.638 0.786
Table 5: Quantitative results of ablation study on each linear edit strength $\lambda$ used for linear edits.
[…]
10 0.431 0.999 0.687 0.924 0.742 0.705 0.646
20 0.568 0.999 0.674 0.921 0.692 0.666 0.724
30 0.649 1.000 0.638 0.900 0.676 0.647 0.755
40 0.708 1.000 0.612 0.890 0.657 0.630 0.773
50 0.739 1.000 0.620 0.883 0.646 0.638 0.786
60 0.771 0.999 0.622 0.892 0.624 0.603 0.791
Table 6: Quantitative results of ablation study on the number of optimization steps.
@@
Ours (tangent) 2.39 1.58 0.80 3.07 1.91 0.99
Diffprivate 3.24 3.52 0.99 3.49 3.24 0.92
Repaint 2.89 1.55 0.97 4.25 2.00 1.22
Diff-Privacy 2.77 1.85 1.33 3.16 1.93 1.57
DeepPrivacy2 3.26 2.71 2.01 3.94 3.09 2.26
CIAGAN 3.44 1.95 1.35 3.69 2.40 1.41
AMT‑GAN 1.51 0.87 0.61 1.15 0.83 0.62
DiffAM 1.50 0.90 0.87 1.57 0.97 0.81
Table 7: Raw pitch, yaw, and roll angle deviations of each method measured by the pose estimator from FaceXFormer. Pitch angle difference from Repaint under FFHQ shows the greatest deviation of $4.25$.
CelebA-HQ FFHQ
Method $downarrow Pitch\downarrow Yaw\downarrow$
Ours (linear) 5.96 4.83 8.45 5.98
Ours (tangent) 5.71 4.72 8.07 5.59
Diffprivate 6.51 4.81 8.00 5.69
Repaint 8.21 6.22 10.66 8.08
Diff-Privacy 10.32 7.16 11.95 8.07
DeepPrivacy2 8.15 5.73 8.66 6.87
CIAGAN 14.45 10.18 14.62 9.98
AMT‑GAN 5.24 5.03 5.55 5.33
DiffAM 9.12 4.79 9.68 5.58
Table 8: Raw pitch and yaw angle deviations of each method measured by L2CS-Net. Pitch angle difference from CIAGAN under FFHQ shows the greatest distortion of $14.62$.
Figure 8: The GUI interface for our user study. Users are asked to decide whether the given images are processed by a de-identification method or not. The image name and source are hidden for fair blind test.
Figure 9: Human recognition of de-identification results based on our user study. FLUID and DeepPrivacy2 are considered the most realistic, showing similar recognition rates to that of CelebA-HQ.
Figure 10: Example images corresponding to each frequently observed user response in our study. Participants frequently recognized low-quality outputs from perturbation-based methods and identified borderline artifacts introduced by inpainting techniques. Additionally, unusual mixtures of gender representations raised user suspicion regarding the authenticity of the data. Notably, the overly perfect appearance of some faces, such as unnaturally clear skin textures and overly smooth hairlines, was also cited as an indicator that the images had been generated or modified by AI models.
Figure 11: Failure examples and biases from out-of-domain FFHQ edit results.
Figure 12: Additional examples demonstrating customizable de-identification. Images are de-identified with the attribute values in the source attribute distribution for ‘Smile’ and ‘Mouth Slightly Open’, and ‘Male’ and ‘Narrow eyes’ fixed from 0.1 to 0.9.
Optimization steps.
In Table 6, we show the impact of the number of optimization steps on the overall performance. The identity-attribute trade-off is also shown here by 10 and 20 steps with low SID and high attribute preservation, indicating insufficiency in de-identification strength and, ultimately, optimization steps. We select 50 steps for our experiments due to the high SID at the cost of a minor decrease in attribute preservation rates. While a higher number of steps yields better overall balance and more iterations might further marginally improve, our results show that 50 steps already achieve competitive performance compared to SOTA methods, validating our approach.
Evaluation thresholds.
To evaluate pose and gaze preservation, we set the thresholds $\tau=5$ and $\tau=15$ for pose and gaze, respectively. These values represent the maximum allowable deviations, beyond which pose and gaze angle differences are regarded as significantly distorted. The thresholds are determined based on the distribution of raw angular deviations, as presented in Tables 7 and 8. In Table 7, Repaint from FFHQ exhibits the largest distortion in pose pitch angle, motivating the choice of $\tau=5$ for pose. Similarly, Table 8 shows that CIAGAN yields the greatest gaze pitch deviation on the FFHQ dataset, justifying the selection of $\tau=15$ for gaze.
Evaluating Human Recognition on De-identified Faces
We conducted a user study to show how human perception detects de-identified faces. We design a simple graphical user interface (GUI) built on Python that shows an image once at a time and asks users to guess whether the image has been de-identified by a generative model or not. We first select $15$ de-identified images each from all compared methods and FLUID, and $20$ real images from CelebA-HQ, with all images chosen from those with high SER-FIQ (terhorst2020ser) scores. This is to prevent the users from identifying fake images from mere artifacts generated by models, which could easily guide the users to identify the fakeness without looking at the faces. We ask $25$ graduate students with diverse nationalities who are not familiar with the field of face de-identification. Note that we do not disclose the number of images selected from each category to the users, nor do we reveal any additional information about the face image that appears in the GUI. An example screenshot of our GUI is shown in Figure 8, while the results and example images are illustrated in Figure 9 and 10, respectively.
Besides FLUID achieving the lowest human recognition rate (0.31), counterintuitively, DeepPrivacy2 has attained a low human recognition rate as well (0.33) despite its low overall score in Table 1. We attribute this to the following responses collected after the blind test, in which the participants were asked to explain the primary factors that led them to conclude an image was de-identified.
@@
Noise-like artifacts from CIAGAN and Diff-Privacy make it obvious to recognize the presense of de-identification. DiffAM’s makeups also belong to this category, offering better results than noise- or patch-based methods, but still suboptimal to be realistic. Users were also able to detect subtle boundaries marked by abrupt color changes across the line, which are commonly observed in mask-based inpainting models such as Repaint. Although we intentionally selected the best-quality results throughout all methods, the typical generation performances of these models remain in the selection.
@@
Customizable de-identification.

### v2

@@ 5.3. Evaluation Metrics
where $F_{id}(\cdot)$ denotes the face identity encoder.
Figure 2. Qualitative comparison with SOTA face de-identification methods on in-domain (CelebA-HQ) and out-of-domain (FFHQ) images. FLUID produces de-identified face images with stricter attribute preservation than other methods. While CelebA-HQ aligns with the pretrained DM’s distribution, FFHQ highlights FLUID’s generalization ability to more diverse real-world inputs.
Attribute Preservation.
To evaluate semantic consistency for realistic de-identification, we focus on gender and age. Gender preservation is estmated using DeepFace (serengil2021hyperextended) as:
(18) $DeepFace(x^)],R_{\text{gender}}=\mathbb{I}\left$$\text$$\text{DeepFace}(\hat{x})\right$],
where $\mathbb{I}[\cdot]$ is the indicator function. The number of results with gender preserved is counted as gender preservation rate.
Age preservation is measured by FaceXFormer’s age estimation function. Specifically, each age range is categorized by age buckets, and we count the age preservation rate just like Equation 18 as:
@@
where $\text{FX}_{age}$ stands for FaceXFormer’s age estimator.
Often, face image generative models distort the poses to alter the perceived geometry and perform de-identification more easily. We additionally evaluate pose consistency using the pose estimator function from FaceXFormer and compute a deviation-based preservation score that combines pitch, yaw, and roll angular differences into a single metric. Given pose deviations $\Delta_{\text{pitch}}$, $\Delta_{\text{yaw}}$, and $\Delta_{\text{roll}}$ between original and de-identified images, we first calculate the root mean square (RMS):
(20) $\text{RMS}_{\text{pose}}=\sqrt{\frac{\Delta_{\text{pitch}}^{2}+\Delta_{\text{yaw}}^{2}+\Delta_{\text{roll}}^{2}}{3}}.$
The pose preservation score is then normalized as:
(21) $R_{\text{pose}}=\max\left(0,1-\frac{\text{RMS}_{\text{pose}}}{\tau}\right),$
where $\tau=5$ represents the angle threshold in degrees for significant pose deviation. Higher scores indicate better pose preservation, with perfect preservation yielding $R_{\text{pose}}=1$. To the end, our results show strong preservation of gender and age for realistic face de-identification, and notably, pose is also well preserved despite the absence of pose-specific supervision.
@@
FALCO 0.499 1.000 0.923 0.658 0.542 0.668 - - - - - -
DeepPrivacy2 0.730 1.000 0.897 0.645 0.458 0.753 0.680 1.000 0.802 0.610 0.366 0.691
Repaint 0.764 0.999 0.871 0.623 0.605 0.794 0.748 0.998 0.721 0.470 0.440 0.703
Diffprivate 0.845 1.000 0.752 0.524 0.436 0.745 0.812 1.000 0.579 0.383 0.440 0.677
FAMS 0.856 1.000 0.431 0.480 0.695 0.729 0.812 1.000 0.642 0.496 0.645 0.761
FLUID (FaceNet, tangent) 0.802 1.000 0.907 0.686 0.674 0.835 0.756 1.000 0.827 0.590 0.584 0.777
FLUID (FaceNet, linear) 0.845 1.000 0.883 0.690 0.646 0.842 0.802 1.000 0.799 0.596 0.562 0.786
FLUID (MagFace, tangent) 0.825 1.000 0.918 0.650 0.620 0.827 0.801 1.000 0.817 0.574 0.592 0.789
FLUID (MagFace, linear) 0.863 1.000 0.889 0.631 0.599 0.829 0.841 0.999 0.801 0.594 0.506 0.784
Table 1. Quantitative performance comparison between FLUID and SOTA face de-identification methods across five key metrics. These include identity removal (SID), image quality (detect), and preservation of three facial attributes: gender, age, and pose. FLUID achieves improved balance between attribute preservation and de-identification, further enhancing realism beyond other SOTA methods. FLUID’s results based on MagFace are described as a replacement analysis under Section 6.4.
Detectability.
When evaluating the quality of face images, FID (heusel2017gans), FIQ (terhorst2020ser), and detectability (king2009dlib; zhang2016joint) are commonly used metrics. FID measures the similarity between the distributions of generated and real data (maximov2020ciagan; li2023riddle; liu2024adv), whereas FIQ quantifies the quality of a face image as perceived by face recognition (FR) models (cai2024disguise). Detectability is often computed as the proportion of images successfully detected by face detection models and is treated as a utility to preserve in anonymization systems (wen2022identitydp; maximov2020ciagan; li2023riddle). However, as FID is based on an Inception network (szegedy2015going) that is trained on ImageNet (deng2009imagenet) without face images, it has recently been considered inappropriate for assessing face image quality (kim2025visual). For FIQ, although it is intended to measure face image quality, it largely captures how easily an image can be processed by face recognition (FR) models, rather than how realistic it appears to human observers. As a result, its scores frequently diverge from intuitive notions of visual fidelity. Therefore, to assess visual realism and face detectability, we report detectability scores as our metric against FR models and complement the results with a user study for human-perceived quality. The number of face detection successes $R_{\text{det}}$ is measured by both Dlib (king2009dlib) and MTCNN (zhang2016joint), and we provide the average detection counts.
Overall Metric.
Evaluating face de-identification systems requires more than checking whether identity is removed; the method must also preserve facial utility and semantic consistency. Existing metrics, however, often reward models that excel in one dimension while quietly failing in others. To address this imbalance, we introduce a nested harmonic mean that captures the overall identity–utility trade-off in a principled way. This is to prevent an aggregate from over‑valuing a configuration that inflates one attribute while degrading another. First, we form a harmonic mean score of three attributes:
(22) $\mathrm{HM}_{\text{att}}=\frac{3}{\frac{1}{R_{\text{gender}}}+\frac{1}{R_{\text{age}}}+\frac{1}{R_{\text{pose}}}},$
then apply another harmonic mean with $R_{\text{id}}$, $R_{\text{det}}$, and $\mathrm{HM}_{\text{att}}$ to obtain the overall score $\mathrm{HM}_{\text{overall}}$. Because harmonic mean is dominated by low components, a method cannot mask weak privacy or a degraded attribute with a single strong metric; nesting enforces internal balance among attributes before combining the three pillars (privacy, detectability, semantics) concisely.
Figure 3. Identity-attribute trade-off curve across increasing linear edit strengths.
6. Results
Quantitative and qualitative comparison.
The closest prior work to ours is FALCO, as both methods optimize a latent vector in the latent space of a generative model with CelebA-defined attributes. However, various differences exist: First, FALCO performs face image de-identification in the popularly used latent space of StyleGAN2 (karras2020analyzing), while we solely rely on the h-space of an unconditional DM for the first time; Second, FALCO is primarily designed for dataset-level anonymization, with the explicit goal of preserving CelebA attribute distributions to maintain dataset utility, rather than for per-image visual fidelity. This design choice is reflected in Table 1 and Figure 2, where FALCO achieves the lowest SID, exhibiting noticeable background distortion, and damaging pose preservation by generating a new image rather than editing the original one; Finally, FLUID improves both visual fidelity and the identity-utility trade-off, thereby extending face de-identification beyond dataset-level anonymization to a more general and flexible setting. This is evidenced by similar attribute preservation performances and significantly improved SID in Table 1, and notably better background preservation in Figure 2.
On CelebA-HQ, FLUID’s tangent edit yields higher gender preservation (0.907) than other models, alongside strong age (0.686) and pose (0.674) consistency. While linear edits show higher SID (0.845) than that of tangent edits (0.802), they often introduce off-manifold artifacts (first and last row of FLUID Linear in Figure 2) and attribute drifts (gender and age pillars in Table 1), which reduce visual realism and semantic stability. Meanwhile, Diffprivate and FAMS achieve higher SID than most of the compared methods by simply sacrificing attributes, producing artifactless but awkward images. Specifically, we observe that Diffprivate tends to smooth the images, making the results femine and younger (FFHQ results of Diffprivate in Figure 2), and FAMS mutually altering both genders and ages (all results of FAMS in Figure 2). Additionally, Diffprivate often attempted to frontalize the faces. All together, these tendencies are evident in the lowest scores of these two methods for gender, age, and pose across both CelebA-HQ and FFHQ. Similarly, as inpainting-based methods, DeepPrivacy2 and Repaint often distort poses (first rows of CelebA-HQ and FFHQ of DeepPrivacy2 results in Figure 2) and offer noticeable borderline artifacts (second rows of CelebA-HQ and FFHQ of Repaint in Figure 2) around the face region, but provide much higher gender and age preservation scores than Diffprivate and FAMS by relying on the knowledge of the trained datasets.
Identity disentanglement validation.
Tangent vs. linear editing.
Why does tangent editing outperform linear editing, and does it generalize? In high-dimensional latent spaces, linear edits often drift off the data manifold, resulting in degraded outputs. Tangent editing, by contrast, constrains the edit direction to the local tangent plane of $h$, preserving both norm and curvature (as shown in Equations 9 and 12). This geodesic formulation avoids semantic drift, significantly improving attribute preservation, visual quality, and cross-domain generalization, particularly in challenging settings such as emotion and pose changes (Table 1).
Figure 4. The GUI interface for our user study. Users are asked to decide whether the given images are processed by a de-identification method or not. The image name and source are hidden for fair blind test.
Figure 5. Human recognition of de-identification results based on our user study. FLUID and DeepPrivacy2 are considered the most realistic, showing low recognition rates similar to that of CelebA-HQ.
Figure 6. Example images corresponding to each frequently observed user response in our study. Participants frequently recognized low-quality outputs from perturbation-based methods and identified borderline artifacts introduced by inpainting techniques. Additionally, unusual mixtures of gender representations raised user suspicion regarding the authenticity of the data. Notably, the overly perfect appearance of some faces, such as unnaturally clear skin textures and overly smooth hairlines, was also cited as an indicator that the images had been generated or modified by AI models.
Figure 7. Qualitative results of ablation study for loss function combinations. The top two rows display CelebA-HQ samples, while the bottom rows are based on FFHQ.
6.1. Evaluating Human Recognition on De-identified Faces
We conducted a user study to evaluate how human perceive de-identified faces. We design a simple graphical user interface (GUI) in Python that shows an image once at a time and asks users to guess whether the image has been de-identified by a generative model or not. We first selected $15$ de-identified images from each compared methods and FLUID, as well as $20$ real images from CelebA-HQ, with all images chosen from those with high SER-FIQ (terhorst2020ser) scores. This is to prevent the users from identifying fake images from mere artifacts generated by models, which could easily guide the users to identify the fakeness without looking at the faces, and automate the user study process to an extent. We recruited $25$ graduate students with diverse nationalities who are not familiar with the field of face de-identification. Note that we do not disclose the number of images selected from each category to the users, nor do we reveal any additional information about the face image that appears in the GUI. An example screenshot of our GUI is shown in Figure 4, while the results and example images are illustrated in Figure 5 and 6, respectively.
Besides FLUID achieving the lowest human recognition rate (0.38) among the compared models, counterintuitively, DeepPrivacy2 has attained low human recognition rate as well (0.41) despite its low overall score in Table 1. We attribute this to the following responses collected after the blind test, in which the participants were asked to explain the primary factors that led them to conclude an image was de-identified.
[…]
8 steps 0.897 0.999 0.873 0.677 0.610 0.848
12 steps 0.841 1.000 0.843 0.647 0.630 0.827
16 steps 0.820 1.000 0.883 0.690 0.646 0.834
20 steps 0.788 0.999 0.876 0.669 0.681 0.824
24 steps 0.749 1.000 0.870 0.649 0.679 0.806
Table 2. Quantitative results on different numbers of denoising steps.
[…]
100 0.589 1.000 0.915 0.695 0.696 0.746
200 0.658 1.000 0.902 0.685 0.677 0.776
300 0.669 1.000 0.904 0.664 0.671 0.777
400 0.683 0.999 0.893 0.681 0.668 0.784
500 0.699 1.000 0.899 0.666 0.665 0.789
600 0.736 1.000 0.880 0.669 0.646 0.800
700 0.765 1.000 0.872 0.665 0.627 0.806
800 0.790 1.000 0.870 0.671 0.624 0.815
900 0.806 1.000 0.877 0.670 0.640 0.824
1000 0.820 1.000 0.883 0.690 0.646 0.834
Table 3. The impact of linear edit strength $\lambda$ used for linear edits.
[…]
10 0.568 0.999 0.924 0.711 0.742 0.743
20 0.679 0.999 0.921 0.695 0.692 0.790
30 0.747 1.000 0.900 0.690 0.676 0.814
40 0.800 1.000 0.890 0.685 0.657 0.829
50 0.820 1.000 0.883 0.690 0.646 0.834
60 0.836 0.999 0.892 0.686 0.624 0.835
Table 4. The tendency of results based on various numbers of optimization steps.
CelebA-HQ FFHQ
[…]
FALCO 2.50 2.85 1.15 - - -
DeepPrivacy2 3.26 2.71 2.01 3.94 3.09 2.26
Repaint 2.89 1.55 0.97 4.25 2.00 1.22
Diffprivate 3.24 3.52 0.99 3.49 3.24 0.92
FAMS 2.13 1.25 0.93 2.25 1.76 1.12
Ours (tangent) 2.39 1.58 0.80 3.07 1.91 0.99
@@ 6.1. Evaluating Human Recognition on De-identified Faces
Table 5. Raw pitch, yaw, and roll angle deviations of each method measured by the pose estimator from FaceXFormer. Pitch angle difference from Repaint under FFHQ shows the greatest deviation of $4.25$.
@@
Artifacts from various methods, though trivial in size and shape, encouraged users to determine such images as fake without looking the face region. Notably, users were also able to detect subtle boundaries marked by abrupt color changes across the line, which are commonly observed in mask-based inpainting models such as Repaint and DeepPrivacy2. Beyond such explicit artifacts, users also flagged images with slight blurriness or monotonous background coloration as fake, which the artifacts are frequently produced by FALCO. De-identification methods operating in latent spaces may incompletely transform or reassign original pixel information (e.g., hat to hair, or hair to background), resulting in residual inconsistencies. Many of the detected results from both FLUID and FALCO fall into this category. Consequently, although we intentionally and automatically selected the highest-quality samples across all methods to minimize such cases, typical generation artifacts remained present in the final selection. This observation further indicates that FIQ is not an ideal metric for assessing face image quality from a perceptual perspective, as it fails to filter out artifact-containing images, particularly those outside the facial region.
@@
6.2. Ablation Study on Loss Functions
To evaluate the individual contribution of the proposed reagent loss components, we conduct an ablation study, where we show qualitative examples in Figure 7. Removing $\mathcal{L}_{\text{ID}}$ removes the objective to displace the original identity, thereby obviously leading to mere reconstruction with trivial changes. Excluding $\mathcal{L}_{\text{att}}$ leads to significant semantic drift, particularly altering key attributes such as gender (top two rows) and facial expression (last row), violating attribute preservation. Other attributes were also discovered to be shifting, such as in the third row where the result without $\mathcal{L}_{\text{att}}$ attempted to add a hat on top of the face. When $\mathcal{L}_{\text{mask}}$ is omitted, we observe varying levels of background artifacts and reduced visual realism. For example, the first and last rows introduce severe artifacts on the background and neck, respectively, while the second and third rows display examples of small portion artifacts on the bottom left region of the images.
Figure 8. Comparison of linear against tangent edits in the h-space. We fit a 3-dimensional principal component analysis (PCA) pool with $5000$ vectors sampled from $100$ editing trajectories, and project the desired vectors onto this PCA space to visualize their trajectory. (a) Linear edits apply raw $\Delta h_{\text{id}}$, allowing the vector norm to drift beyond the typical shell and producing visual artifacts. (b) Tangent edits project $\Delta h_{\text{id}}$ onto the local tangent plane at a constant norm, constraining the trajectory and reducing artifacts.
6.3. Influence of Hyperparameter Selection and Analysis
Denoising steps and computational analysis.
Similarly, we conduct another comparison to select the number of denoising steps. As written in Table 2, increasing the number of reverse steps tends to reduce SID and retain attributes, especially pose. Although 8 denoising steps show the best balance due to the strongest SID, we select 16 steps for our experiments for improved stability and consistency, ensuring reliable performance across diverse inputs and more attribute preservation. Even though the number of denoising steps is much less than that of DDPM thanks to the implementation of DDIM (song2020denoising) inversion and generation, naively backpropagating through the iterative reverse steps still requires considerable amount of memory. Therefore, we implement gradient checkpointing (chen2016training) to significantly reduce memory consumption, enabling us to run the implementation on our hardware, a single Titan XP GPU with 12GB of memory. With the help of gradient checkpointing, FLUID runs with approximately 6GB of memory and takes under 3 seconds for each optimization step, including all auxiliary model operations.
(23) $p_{\theta}^{(t)}(x_{t-1}\mid x_{t})=\mathcal{N}\left(\sqrt{\alpha_{t-1}}\,\mathbf{P}_{t}\left(\epsilon_{\theta}^{t}(x_{t})\right)+\mathbf{D}_{t},\sigma_{t}^{2}\mathbf{I}\right).$
Linear edit strength.
We clarify that the editing strength $\lambda$ for linear edits is set to $1000$. In contrast, Boundary Diffusion offers the editing strength within the range of [150,150][-150,150], as its direction vector encodes the signed distance to a hyperplane (i.e., boundary) defined by a linear SVM. Since we do not rely on such boundaries during optimization, and the norm of $\Delta h_{\text{id}}$ varies throughout the process, we conduct a comparison study on the linear editing strength $\lambda$ in the range of $100$ to $1000$, as shown in Table 3. Unlike other replacement experiments where trade-offs are evident, we observe trivial trade-off patterns in this case. Larger values of $\lambda$ consistently lead to increased SID, with constantly high detectability scores and slightly decreasing attribute preservation, thereby improving the overall score. We attribute the increase in SID to a larger exploration range enabled by stronger scaling factors within a limited number of optimization steps. Based on this observation, we select $\lambda=1000$ as our default linear edit strength.
Figure 9. Pareto front curves and data points plotted for each dataset. The Pareto front is formed constantly with FLUID’s data points, while other models are located behind the front.
Optimization steps.
In Table 4, we show the impact of the number of optimization steps on the overall performance. The identity-attribute trade-off is also shown here by 10 and 20 steps with low SID and high attribute preservation, indicating insufficiency in de-identification strength and, ultimately, optimization steps. Although 30 steps already outperforms other SOTA methods, we select 50 steps for our experiments due to the high SID at the cost of a minor decrease in attribute preservation rates. While a higher number of steps yields better overall balance and more iterations might further marginally improve, our results show that 50 steps already achieve competitive performance compared to SOTA methods, while results around 40 steps start to plateau.
6.4. The FR Model’s Impact for White-box Attack
We provide an example when the FR model that is used for white-box attacks is altered. For this purpose, selecting a fundamentally different FR model from FaceNet will yield meaningful outcomes. We replace FaceNet, an L2 distance-based model, with MagFace (meng2021magface), an angular distance-based FR model where cosine similarity correlates with identity similarity. Since FaceNet and MagFace differs in design (i.e., the distance metric), we introduce a new identity loss function suitable for MagFace. Given a source image $x$ and a generated image $\hat{x}$, we define this identity loss as:
[…]
where $F_{x}$ denotes the MagFace embedding of image $x$ and CS stands for cosine similarity. In Table 1, we append MagFace’s performance after that of FaceNet. The improvement over utilities is also shown by MagFace-based linear and tangent edits, aligning with our hypothesis. It is important to note that white-box attacks against MagFace produce the highest SID values and yet higher balance scores compared to other SOTA baselines. However, the age and pose scores are sacrificed compared to FaceNet results, posing the identity-utility trade-off.
Figure 10. Additional examples demonstrating customizable de-identification. Images are de-identified with the attribute values in the source attribute distribution for ‘Smile’ and ‘Mouth Slightly Open’, and ‘Male’ and ‘Narrow eyes’ fixed from 0.1 to 0.9.
6.5. Visualizing the Identity-Utility Trade-off
Although the privacy-utility trade-off is quantitatively measured by offering a nested harmonic mean, we additionally visualize this trade-off by presenting a Pareto front curve for a more comprehensive and fair comparison. We select two pillars from Table 1; SID and the harmonic mean of three attributes. We do not include detectability as the third pillar because the difference is trivial across all models (up to $0.2\%$). The Pareto front curve highlights models that achieve an optimal balance between these competing objectives, in the sense that no model on the curve can improve one pillar without degrading the other. Models lying below the curve are strictly dominated, indicating inferior trade-offs compared to those on the Pareto frontier. As displayed in Figure 9, FaceNet-based FLUID and the linear variant of MagFace form a Pareto front in CelebA-HQ, while all FLUID results are included in the Pareto front, highlighting the superior balance between identity removal and utility preservation quantitatively. Meanwhile, the balance of FAMS differs significantly across two datasets due to the lowest gender and age preservation scores at CelebA-HQ. FALCO and DeepPrivacy2 locate far from the Pareto front, which is supported by the lowest SID scores.
6.6. Customizable de-identification.